# 1in1in1in1in

### v2

@@ 3.2 Property $\mathbf{(b)}$: Homeomorphism on $\mathbf{\mathbb{P}T_{p}S}$
By postcomposing with a rotation, we can assume that $h$ fixes some point $c\in S^{1}$. Let $\varphi:S^{1}\setminus\{c\}\rightarrow\mathbb{R}$ be a homeomorphism such that for any $x,y\in\mathbb{R}$ such that $x<y$, the cyclic triple $[\varphi^{-1}(x)\varphi^{-1}(y)c]$ of $S^{1}$ is in clockwise order.
Refer to caption $\varphi$$\varphi^{-1}(x)$$\varphi^{-1}(y)$$x$$y$
Figure 9: The homeomorphism $\varphi$ from $S^{1}\setminus\{c\}$ to $\mathbb{R}$.
Then $\varphi\circ h|_{S^{1}\setminus\{c\}}\circ\varphi^{-1}$ is an order preserving (or reversing) bijection on $\mathbb{R}$. So by Observation 3.9, $\varphi\circ h|_{S^{1}\setminus\{c\}}\circ\varphi^{-1}$ is a homeomorphism. Thus $h|_{S^{1}\setminus\{c\}}$ is also a homeomorphism. Moreover, since any neighborhood of $c$ will be sent by $h|_{S^{1}\setminus\{c\}}$ to a neighborhood of $c$, then $h|_{S^{1}\setminus\{c\}}$ can be extended to $h$ and remain a homeomorphism on all of $S^{1}$. ∎
@@
Let $[abc]$ be a cyclic triple of $\mathbb{P}T_{p}S$ in clockwise order. Since $a$, $b$, and $c$ are distinct, then there are representative curves $\alpha$, $\beta$, and $\gamma$ that are transverse at $p$ and a connected neighborhood $N$ of $p$ such that these curves are disjoint in $N\setminus\{p\}$. Denote these disjoint arcs as $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$, $\gamma_{1}$, and $\gamma_{2}$ such that in clockwise order around $p$, these arcs occur as the cycle $[\alpha_{1}\beta_{1}\gamma_{1}\alpha_{2}\beta_{2}\gamma_{2}]$.
Refer to caption $f$$\bar{d}f_{p}$$a$$b$$c$$\alpha_{1}$$\alpha_{2}$$\beta_{1}$$\beta_{2}$$\gamma_{1}$$\gamma_{2}$
Figure 10: The correspondence between cyclic triples in $\mathbb{P}T_{p}S$ and their representatives on $S$ near the point $p$.
Since $f$ is a homeomorphism, then the arcs $f(\alpha_{1}),f(\alpha_{2}),f(\beta_{1}),f(\beta_{2}),(\gamma_{1})$, and $f(\gamma_{2})$ are also disjoint in $f(N\setminus\{p\})$. Moreover, the cycle $[f(\alpha_{1})f(\beta_{1})f(\gamma_{1})f(\alpha_{2})f(\beta_{2})f(\gamma_{2})]$ is in clockwise order since $f$ is orientation preserving. Thus $f(\alpha)$, $f(\beta)$, and $f(\gamma)$ are still topologically transverse at $p$ with the same ordering.
By Lemma 3.8, we know that $\bar{d}f_{p}$ is a bijection. Thus $\bar{d}f_{p}(a),\bar{d}f_{p}(b),$ and $\bar{d}f_{p}(c)$ are distinct and the cyclic triple $[\bar{d}f_{p}(a)\,\bar{d}f_{p}(b)\,\bar{d}f_{p}(c)]$ is in clockwise order. Thus $\bar{d}f_{p}$ preserves the cyclic order of every triple. ∎
@@
By Lemmas 3.8 and 3.11, we have that for any $f\in\operatorname{Homeo}^{1}(S)$, the map $\bar{d}f_{p}$ satisfies the conditions of Lemma 3.10. Thus $\bar{d}f_{p}$ is a homeomorphism for every $p\in S$. ∎
@@ 3.3 Property $\mathbf{(c)}$: Mapping transverse sequences
We complete this proof in two steps. We first find the subsequence of points $\{p_{n_{k}}\}$ and then we find the $C^{1}$ curve $\gamma$ that passes through this subsequence. This construction is inspired by a similar proof from Rosenthal [Ros55, Theorem 2] that finds a $C^{1}$ curve through a bounded sequence of points. We have made adjustments to their argument to account for the additional restriction on tangent lines.
@@
$\gamma_{k}(t)$ =$=\left\{\begin{array}[]{ll}\left(\frac{2t-x_{k}-x_{k+1}}{x_{k}-x_{k+1}}\right)\left(\left(\frac{(x_{k}-x_{k+1})m_{k}}{2\pi}\right)\,\sin\left(\frac{2\pi(t-x_{k})}{x_{k}-x_{k+1}}\right)+y_{k}\right)&\\ \qquad+\left(1-\left(\frac{2t-x_{k}-x_{k+1}}{x_{k}-x_{k+1}}\right)\right)\left(\frac{y_{k}-y_{k+1}}{x_{k}-x_{k+1}}(t-x_{k})+y_{k}\right)&\mbox{ if }\frac{x_{k}+x_{k+1}}{2}\leq t\leq x_{k}\\ \left(\frac{2t-x_{k}-x_{k+1}}{x_{k+1}-x_{k}}\right)\left(\left(\frac{(x_{k}-x_{k+1})m_{k+1}}{2\pi}\right)\,\sin\left(\frac{2\pi(t-x_{k+1})}{x_{k}-x_{k+1}}\right)+y_{k+1}\right)&\\ \qquad+\left(1-\left(\frac{2t-x_{k}-x_{k+1}}{x_{k+1}-x_{k}}\right)\right)\left(\frac{y_{k}-y_{k+1}}{x_{k}-x_{k+1}}(t-x_{k+1})+y_{k+1}\right)&\mbox{ if }x_{k+1}\leq t\leq\frac{x_{k}+x_{k+1}}{2}\end{array}\right.$
@@
$\gamma_{k}^{\prime}(t)=\left\{\begin{array}[]{ll}m_{k}\left(\left(\frac{2t-x_{k}-x_{k+1}}{x_{k}-x_{k+1}}\right)\cos\left(\frac{2\pi(t-x_{k})}{x_{k}-x_{k+1}}\right)+\frac{1}{\pi}\sin\left(\frac{2\pi(t-x_{k})}{x_{k}-x_{k+1}}\right)\right)&\\ \qquad+\;2\left(1-\frac{2t-x_{k}-x_{k+1}}{x_{k}-x_{k+1}}\right)\left(\frac{y_{k}-y_{k+1}}{x_{k}-x_{k+1}}\right)&\mbox{ if }\frac{x_{k}+x_{k+1}}{2}\leq t\leq x_{k}\\ m_{k+1}\left(\left(\frac{2t-x_{k}-x_{k+1}}{x_{k+1}-x_{k}}\right)\cos\left(\frac{2\pi(t-x_{k+1})}{x_{k}-x_{k+1}}\right)-\frac{1}{\pi}\sin\left(\frac{2\pi(t-x_{k+1})}{x_{k}-x_{k+1}}\right)\right)&\\ \qquad+\;2\left(1-\frac{2t-x_{k}-x_{k+1}}{x_{k+1}-x_{k}}\right)\left(\frac{y_{k}-y_{k+1}}{x_{k}-x_{k+1}}\right)&\mbox{ if }x_{k+1}\leq t\leq\frac{x_{k}+x_{k+1}}{2}\\ \end{array}\right.$
@@
Refer to caption $\alpha$$\xi$$\beta$$\delta$
Figure 11: The closed $C^{1}$ curve $\gamma$ in $\mathbb{R}^{2}$.
@@
Refer to caption $\ell$$\gamma^{+}$$\gamma^{-}$$-\varepsilon$$\varepsilon$
Figure 12: The points $(x_{n},y_{n})$ converging along the line $\ell$ eventually stay between the regions defined by $(-\varepsilon-\delta/2,-\delta/2)\cup(\delta/2,\varepsilon+\delta/2)$ and thus also the curves $\gamma^{+}$ and $\gamma^{-}$.
@@
We now prove the reverse direction. Let $\gamma^{+}$ and $\gamma^{-}$ be the straight lines with angles $\varepsilon$ and $-\varepsilon$. Since $\varepsilon$ is arbitrary, then the angle of the points $(x_{n},y_{n})$ relative to the origin must be converging to 0. Thus the sequence $\{(x_{n},y_{n})\}$ is converging along the $x$-axis to the origin. ∎
Note that Lemma 3.15 also applies locally to sequences that converge along lines to points in a surface $S$ by replacing straight line representatives with locally geodesic representatives. We can now prove Lemma 3.12 and Propostion 3.1, giving the forward direction of our Main Theorem.
@@
Proof of Proposition 3.1.
Property $(a)$ comes directly from the definition of $\operatorname{Homeo^{1}}(S)$. The other two desired properties are given by Lemmas 3.7 and 3.12, respectively. ∎
@@ 4.1 Well-defined tangent lines
But the limit for all the secant slopes approaching 00 does not exist, so there is some open neighborhood $N_{m}$ of $m$ such that there exist $s_{n}\in\left(-\frac{1}{n},\frac{1}{n}\right)$ with
@@
By Observation 4.3, there exist simple closed $C^{1}$ curves $\alpha$ and $\beta$ such that $\alpha$ intersects infinitely many of the $f^{-1}(\gamma(t_{n}))$ and $\beta$ intersects infinitely many of the $f^{-1}(\gamma(s_{n_{k}}))$. Note that the tangent line of $\alpha$ at $p$ must have a slope of $m$, while the tangent line of $\beta$ at $p$ must have a slope of $m^{\prime}$. Since $f(\alpha)$ and $\gamma$ are $C^{1}$ curves that intersect at infinitely many distinct points that converge to $f(p)$, then by Corollary 3.6, $f(\alpha)$ and $\gamma$ must have the same tangent line at $f(p)$. Similarly, $f(\beta)$ and $\gamma$ are $C^{1}$ curves with the same tangent line at $f(p)$. Thus $f(\alpha)$ and $f(\beta)$ must have the same tangent line at $f(p)$. But since $\bar{d}f_{p}$ is a homeomorphism, this means that $\alpha$ and $\beta$ must have the same tangent line at $p$. This is a contradiction since by construction $m\neq m^{\prime}$. ∎
@@ 4.2 Continuity of tangent lines
Every $C^{1}$ curve has tangent lines at every point that vary continuously. Work of Burgués–Cufí [BC09, Theorem 10] shows that for simple closed curves, continuously varying tangent lines is sufficient to recover a $C^{1}$-regular parameterization with nonzero tangent vectors.
@@ 3. $(c)$
Now suppose for a contradiction that $M_{f^{-1}(\gamma)}$ is not continuous. Then there is a sequence of $t_{n}$ in $\mathbb{R}$ that converges to 00 such that $M_{f^{-1}(\gamma)}(t_{n})$ does not converge to $M_{f^{-1}(\gamma)}(0)$. Moreover, there exists an $\varepsilon>0$ and a subsequence $t_{n_{k}}$ such that $M_{f^{-1}(\gamma)}(t_{n_{k}})$ is outside of an $\varepsilon$-neighborhood of $M_{f^{-1}(\gamma)}(0)$. Thus no subsequence of $M_{f^{-1}(\gamma)}(t_{n_{k}})$ converges to $M_{f^{-1}(\gamma)}(0)$. Let $p_{k}$ denote the points $f^{-1}(\gamma)(t_{n_{k}})$ and let $\ell_{k}$ denote the element in $\mathbb{P}T_{p_{k}}S$ that corresponds to $M_{f^{-1}(\gamma)}(t_{n_{k}})$. So $\{(p_{k},\ell_{k})\}$ is a transverse sequence.
@@
Proof of Proposition 4.1.
By Lemmas 4.2 and 4.4, $f^{-1}$ maps a $C^{1}$ curve $\gamma$ to a curve that has continuous tangent line at every point. The result of Burgués–Cufí [BC09, Theorem 10] implies that $f^{-1}(\gamma)$ is also a $C^{1}$ curve. ∎
@@ 4.3 Finishing the proof
Combining Propositions 3.1 and 4.1, we now derive our main result.
@@
Proposition 3.1 proves the forward direction of the theorem. For the reverse direction, property $(a)$ gives that $f$ maps every $C^{1}$ curve to a $C^{1}$ curve. By Proposition 4.1, $f^{-1}$ also maps every $C^{1}$ curve to a $C^{1}$ curve. Thus $f\in\operatorname{Homeo^{1}}(S)$. ∎

### v3

@@ 3.2 Property $\mathbf{(b)}$: Homeomorphism on $\mathbf{\mathbb{P}T_{p}S}$
By postcomposing with a rotation, we can assume that $h$ fixes some point $c\in S^{1}$. Let $\varphi:S^{1}\setminus\{c\}\rightarrow\mathbb{R}$ be a homeomorphism such that for any $x,y\in\mathbb{R}$ such that $x<y$, the cyclic triple $\varphi^{-1}(x)\varphi^{-1}(y)c]$ of $S^{1}$ is in clockwise order.
Figure 9: The homeomorphism $\varphi$ from $S^{1}\setminus\{c\}$ to $\mathbb{R}$
Then $\varphi\circ h|_{S^{1}\setminus\{c\}}\circ\varphi^{-1}$ is an order preserving (or reversing) bijection on $\mathbb{R}$. So by Observation 3.8, $\varphi\circ h|_{S^{1}\setminus\{c\}}\circ\varphi^{-1}$ is a homeomorphism. Thus $h|_{S^{1}\setminus\{c\}}$ is also a homeomorphism. Moreover, since any neighborhood of $c$ will be sent by $h|_{S^{1}\setminus\{c\}}$ to a neighborhood of $c$, then $h|_{S^{1}\setminus\{c\}}$ can be extended to $h$ and remain a homeomorphism on all of $S^{1}$. ∎
@@
Let $[abc]$ be a cyclic triple of $\mathbb{P}T_{p}S$ in clockwise order. Since $a$, $b$, and $c$ are distinct, then there are representative curves $\alpha$, $\beta$, and $\gamma$ that are transverse at $p$ and a connected neighborhood $N$ of $p$ such that these curves are disjoint in $N\setminus\{p\}$. Denote these disjoint arcs as $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$, $\gamma_{1}$, and $\gamma_{2}$ such that in clockwise order around $p$, these arcs occur as the cycle $\alpha_{1}\beta_{1}\gamma_{1}\alpha_{2}\beta_{2}\gamma_{2}]$.
Figure 10: The correspondence between cyclic triples in $\mathbb{P}T_{p}S$ and their representatives on $S$ near the point $p$
Since $f$ is a homeomorphism, then the arcs $f(\alpha_{1}),f(\alpha_{2}),f(\beta_{1}),f(\beta_{2}),(\gamma_{1})$, and $f(\gamma_{2})$ are also disjoint in $f(N\setminus\{p\})$. Moreover, the cycle $f(\alpha_{1})f(\beta_{1})f(\gamma_{1})f(\alpha_{2})f(\beta_{2})f(\gamma_{2})]$ is in clockwise order since $f$ is orientation preserving. Thus $f(\alpha)$, $f(\beta)$, and $f(\gamma)$ are still topologically transverse at $p$ with the same ordering.
By Lemma 3.7, we know that $\bar{d}f_{p}$ is a bijection. Thus $\bar{d}f_{p}(a),\bar{d}f_{p}(b),$ and $\bar{d}f_{p}(c)$ are distinct and the cyclic triple $\bar{d}f_{p}(a)\,\bar{d}f_{p}(b)\,\bar{d}f_{p}(c)]$ is in clockwise order. Thus $\bar{d}f_{p}$ preserves the cyclic order of every triple. ∎
@@
By Lemmas 3.7 and 3.10, we have that for any $f\in\operatorname{Homeo}^{1}(S)$, the map $\bar{d}f_{p}$ satisfies the conditions of Lemma 3.9. Thus $\bar{d}f_{p}$ is a homeomorphism for every $p\in S$. ∎
@@ 3.3 Property $\mathbf{(c)}$: Mapping transverse sequences
We complete this proof in two steps. We first find the subsequence of points $\{p_{n_{k}}\}$ and then we find the $C^{1}$ curve $\gamma$ that passes through this subsequence. This construction is inspired by a similar proof from Rosenthal [4, Theorem 2] that finds a $C^{1}$ curve through a bounded sequence of points. We have made adjustments to their argument to account for the additional restriction on tangent lines.
@@
$\gamma_{k}(t)$ $=\left\{\begin{array}[]{ll}\left(\frac{2t-x_{k}-x_{k+1}}{x_{k}-x_{k+1}}\right)\left(\left(\frac{(x_{k}-x_{k+1})m_{k}}{2\pi}\right)\,\sin\left(\frac{2\pi(t-x_{k})}{x_{k}-x_{k+1}}\right)+y_{k}\right)&\\ \qquad+\left(1-\left(\frac{2t-x_{k}-x_{k+1}}{x_{k}-x_{k+1}}\right)\right)\left(\frac{y_{k}-y_{k+1}}{x_{k}-x_{k+1}}(t-x_{k})+y_{k}\right)&\mbox{ if }\frac{x_{k}+x_{k+1}}{2}\leq t\leq x_{k}\\ \left(\frac{2t-x_{k}-x_{k+1}}{x_{k+1}-x_{k}}\right)\left(\left(\frac{(x_{k}-x_{k+1})m_{k+1}}{2\pi}\right)\,\sin\left(\frac{2\pi(t-x_{k+1})}{x_{k}-x_{k+1}}\right)+y_{k+1}\right)&\\ \qquad+\left(1-\left(\frac{2t-x_{k}-x_{k+1}}{x_{k+1}-x_{k}}\right)\right)\left(\frac{y_{k}-y_{k+1}}{x_{k}-x_{k+1}}(t-x_{k+1})+y_{k+1}\right)&\mbox{ if }x_{k+1}\leq t\leq\frac{x_{k}+x_{k+1}}{2}\end{array}\right$.
@@
[…]
@@
Figure 11: The closed $C^{1}$ curve $\gamma$ in $\mathbb{R}^{2}$
@@
Figure 12: The points $(x_{n},y_{n})$ converging along the line $\ell$ eventually stay between the regions defined by $(-\varepsilon-\delta/2,-\delta/2)\cup(\delta/2,\varepsilon+\delta/2)$ and thus also the curves $\gamma^{+}$ and $\gamma^{-}$
@@
We now prove the reverse direction. Let $\gamma^{+}$ and $\gamma^{-}$ be the straight lines with angles $\varepsilon$ and $\varepsilon$. Since $\varepsilon$ is arbitrary, then the angle of the points $(x_{n},y_{n})$ relative to the origin must be converging to 0. Thus the sequence $\{(x_{n},y_{n})\}$ is converging along the $x$-axis to the origin. ∎
Note that Lemma 3.14 also applies locally to sequences that converge along lines to points in a surface $S$ by replacing straight line representatives with locally geodesic representatives. We can now prove Lemma 3.11 and Propostion 8, giving the forward direction of our Main Theorem.
@@
Proof of Proposition 8.
Property $(a)$ comes directly from the definition of $\operatorname{Homeo^{1}}(S)$. The other two desired properties are given by Lemmas 3.6 and 3.11, respectively. ∎
@@ 4.1 Well-defined tangent lines
But the limit for all the secant slopes approaching 0 does not exist, so there is some open neighborhood $N_{m}$ of $m$ such that there exist $s_{n}\in\left(-\frac{1}{n},\frac{1}{n}\right)$ with
@@
By Observation 4.2, there exist simple closed $C^{1}$ curves $\alpha$ and $\beta$ such that $\alpha$ intersects infinitely many of the $f^{-1}(\gamma(t_{n}))$ and $\beta$ intersects infinitely many of the $f^{-1}(\gamma(s_{n_{k}}))$. Note that the tangent line of $\alpha$ at $p$ must have a slope of $m$, while the tangent line of $\beta$ at $p$ must have a slope of $m^{\prime}$. Since $f(\alpha)$ and $\gamma$ are $C^{1}$ curves that intersect at infinitely many distinct points that converge to $f(p)$, then by Corollary 3.5, $f(\alpha)$ and $\gamma$ must have the same tangent line at $f(p)$. Similarly, $f(\beta)$ and $\gamma$ are $C^{1}$ curves with the same tangent line at $f(p)$. Thus $f(\alpha)$ and $f(\beta)$ must have the same tangent line at $f(p)$. But since $\bar{d}f_{p}$ is a homeomorphism, this means that $\alpha$ and $\beta$ must have the same tangent line at $p$. This is a contradiction since by construction $m\neq m^{\prime}$. ∎
@@ 4.2 Continuity of tangent lines
Every $C^{1}$ curve has tangent lines at every point that vary continuously. Work of Burgués–Cufí [5, Theorem 10] shows that for simple closed curves, continuously varying tangent lines is sufficient to recover a $C^{1}$-regular parameterization with nonzero tangent vectors.
@@ 3. $(c)$
Now suppose for a contradiction that $M_{f^{-1}(\gamma)}$ is not continuous. Then there is a sequence of $t_{n}$ in $\mathbb{R}$ that converges to 0 such that $M_{f^{-1}(\gamma)}(t_{n})$ does not converge to $M_{f^{-1}(\gamma)}(0)$. Moreover, there exists an $\varepsilon>0$ and a subsequence $t_{n_{k}}$ such that $M_{f^{-1}(\gamma)}(t_{n_{k}})$ is outside of an $\varepsilon$-neighborhood of $M_{f^{-1}(\gamma)}(0)$. Thus no subsequence of $M_{f^{-1}(\gamma)}(t_{n_{k}})$ converges to $M_{f^{-1}(\gamma)}(0)$. Let $p_{k}$ denote the points $f^{-1}(\gamma)(t_{n_{k}})$ and let $\ell_{k}$ denote the element in $\mathbb{P}T_{p_{k}}S$ that corresponds to $M_{f^{-1}(\gamma)}(t_{n_{k}})$. So $\{(p_{k},\ell_{k})\}$ is a transverse sequence.
@@
Proof of Proposition 9.
By Lemmas 4.1 and 4.3, $f^{-1}$ maps a $C^{1}$ curve $\gamma$ to a curve that has continuous tangent line at every point. The result of Burgués–Cufí [5, Theorem 10] implies that $f^{-1}(\gamma)$ is also a $C^{1}$ curve. ∎
@@ 4.3 Finishing the proof
Combining Propositions 8 and 9, we now derive our main result.
@@
Proposition 8 proves the forward direction of the theorem. For the reverse direction, property $(a)$ gives that $f$ maps every $C^{1}$ curve to a $C^{1}$ curve. By Proposition 9, $f^{-1}$ also maps every $C^{1}$ curve to a $C^{1}$ curve. Thus $f\in\operatorname{Homeo^{1}}(S)$. ∎